# LibMOON: A Gradient-based MultiObjective OptimizatioN Library in PyTorch

### v2

@@ 4 Empirical studies
In this section, we study empirical results of LibMOON. Codes are basically run on a person computer with an i7-10700 CPU, and a 3080 GPU with 10G GPU memory. And we also conduct some analysis regarding to GPU ranging from 3080, 4060, and 4090. The empirical results contain five parts, MOO solvers for synthetic problems, Pareto set learning solvers for synthetic problems, MOO for MTL problems, Pareto set learning solvers for MTL problems, and GPU accelerations. The report of the first subsections .
@@ 4.1 MOO solvers for synthetic problems
We report the performance of various finite Pareto solution solvers using the VLMOP2 problem, formulated as follows:
[…]
This problem has been widely studied in the literature [14, 18] due to its PF has a non-convex shape  [^2] In MOO, a PF is convex or non-convex based on whether the objective space is convex or non-convex. The PF represents the non-dominated boundary of the objective space.. The visualization results are shown in Figure 3 and the numerical results are shown in Table 6.
@@
(k) Random
Figure 3: Finite Pareto solutions by different solvers on VLMOP2 problem.
@@
Method Lmin Smooth Lmin Spacing Sparsity HV IP Cross Angle PBI
@@
  • All methods were run five times with random seeds, presented as (mean)(std).
We present the following key findings:
  1. 1.
    Agg-COSMOS produces approximate ‘exact’ Pareto solutions (the corresponding Pareto objectives aligned with preference vectors). This is due to a cosine similarity term in Agg-COSMOS encouraging the Pareto objectives be close to preference vectors.
  2. 2.
    Agg-LS can only find two endpoints on this problem since PF of VLMOP2 is of non-convex shape. Different preference vectors correspond to duplicated Pareto solutions.
  3. 3.
    Agg-PBI generates “exact” Pareto solutions when the coefficient of $d_{2}$ exceeds a specific value, as guaranteed by PBI [54]. However, this parameter is challenging to tune and is influenced by the Pareto front’s curvature. Additionally, PBI can transform a convex multi-objective optimization problem into a non-convex one, making Agg-PBI less recommended.
  4. 4.
    Agg-Tche generates diverse solutions and produces exact Pareto solutions corresponding to the inverse of the preference vector. Both Agg-Tche and Agg-SmoothTche retain the convexity of objectives, meaning their aggregation functions remain convex when all objectives are convex. However, the solution distribution of Agg-SmoothTche cannot be precisely determined due to the involvement of a temperature coefficient $h$.
  5. 5.
    HVGrad updates the decision variable using the hypervolume gradient, resulting in the largest hypervolume. PMTL, a two-stage method, can struggle with tuning warm-up iterations, leading to poor initialization and less diverse solutions. MOO-SVGD’s update direction has two conflicting goals: promoting diversity and ensuring convergence. This conflict makes MOO-SVGD difficult to converge, often requiring 10 times more iterations than other methods, resulting in poor convergence behavior.
  6. 6.
    Among the methods, MGDA-UB, Random, Agg-PBI, and MOO-SVGD exhibit relatively large deviations. MGDA-UB and Random generate arbitrary Pareto solutions due to their computation nature. Agg-PBI results in non-convex aggregation functions, leading to variable solutions based on different random initializations. MOO-SVGD has high standard deviation because its optimization process is unstable, and its balance between convergence and diversity is unclear.
@@ 4.2 Pareto set learning on synthetic problems
In this section, we present PSL results using the synthetic VLMOP2 problem as an example. Visualizations of the predicted Pareto solutions are shown in Figure 4, and numerical results are reported in Table 7. From the table and figure, we have the following key findings:
@@
Method Lmin Smooth Lmin Spacing Sparsity HV IP Cross Angle PBI Span
Agg-COSMOS 0.045 (0.000) -0.127 (0.000) 1.560 (0.004) 0.525 (0.000) 0.318 (0.000) 0.752 (0.000) 0.950 (0.001) 0.995 (0.000) 0.907 (0.000)
@@
  1. 1.
    PSL with the COSMOS aggregation function fails to find all marginal Pareto solutions because COSMOS does not guarantee the discovery of the entire Pareto set/front. PSL with linear scalarization function could not fit the two endpoints of the PF.
  2. 2.
    PSL with the smooth Tchebycheff function finds diverse Pareto solutions due to the unclear relationship between preference vectors and Pareto objectives. PSL using Tche., EPO, and PMGDA as base solvers can discover the entire PS/PF, as all three methods effectively find exact Pareto solutions. By traversing the entire preference space, the Pareto model fits the entire Pareto set accurately.
(a) Agg-COSMOS
@@ 4.3 MOO solvers for MTL problems
In this section, we investigate the performance of different finite Pareto solution solvers on the Adult dataset, which is a multitask learning task called fairness classification. For this task, the decision variable $\theta$ is the parameter of a fully-connected neural network, with $|{\bm{\theta}}|=28033$. The first objective function is the cross entropy loss, and the second objective is the hyperbolic tangent relaxation of the difference of equality of opportunity (DEO) loss [37][Eq. 6]. The visualization on the Adult problem is shown in Figure 5 and the numerical result is shown in Table 8.
From the table and figure, we have the following key findings:
  1. 1.
    Agg-LS has two main drawbacks: it cannot identify the non-convex parts of the Pareto front (as previously mentioned), and the relationship between preference vectors and Pareto objectives is unknown, making it challenging to select appropriate preference vectors. Agg-PBI and Agg-COSMOS only find a small portion of the PF due to it is hard to set the coefficients.
  2. 2.
    Agg-SmoothmTche and Agg-mTche perform well on this task, as they can accurately find Pareto solutions. Once the Pareto front range is known, diverse solutions can be easily found using uniform preference vectors.
  3. 3.
    The Random method and MGDA-UB only find a single Pareto solution, since the position of this solution cannot be controlled by these methods.
  4. 4.
    Among the three methods that directly find a set of Pareto solutions (MOO-SVGD, PMTL, and HV-Grad), HV-Grad produces the most diverse solutions with the largest hypervolume. PMTL, being a two-stage method, can encounter issues when solutions fall outside the sector due to stochastic factors. MOO-SVGD optimizes both convergence and diversity but is generally unstable based on our tests.
@@
Method Lmin Smooth Lmin Spacing Sparsity HV IP Cross Angle PBI Span
@@
  • All methods were run five times with random seeds, presented as (mean)(std).
@@ 4.4 Pareto set learning on MTL
This section presents the Pareto set learning results on the MO-MNIST problem using a hypernet architecture. The model was trained for 20 epochs, optimizing approximately 3.24M parameters, with the first and second objectives being the cross-entropy losses for the top-right and bottom-left images. EPO-based PSL and PMGDA-based PSL are not very suitable for this task since manipulating gradient of 3.2M dim is not efficient, the visualization results are reported in Figure 6 and numerical results in Table 9.
From the table and figure, we have the following key findings:
  1. 1.
    The Pareto Front (PF) of this task is nearly convex, indicating that Linear scalarazation (LS) can efficiently find the entire Pareto front. Therefore, it is not necessary to employ complex Multi-Objective Optimization (MOO) methods.
  2. 2.
    Agg-LS significantly outperforms other methods on this task, as evidenced by the substantially higher Hypervolume (HV). Furthermore, the training losses across other methods show minimal reduction, indicating that for this challenging task, the convergence of other non-linear scalarization methods is extremely slow.
  3. 3.
    Agg-SmoothTche and Agg-Tche produce duplicate marginal Pareto solutions because selecting appropriate weight factors is challenging.
@@
Method Lmin Smooth Lmin Spacing Sparsity HV IP Cross Angle PBI Span
Agg-COSMOS 0.033 (0.002) -0.152 (0.007) 0.967 (0.202) 0.278 (0.038) 0.512 (0.015) 0.535 (0.028) 9.208 (0.541) 1.221 (0.050) 0.497 (0.043)
@@ 4.5 MOBO for synthetic and real-world problems
4.6 GPU acceleration
Table 10: Running time (minutes) for different platforms on the MO-MNIST problem.
Platform MO-MNIST MO-Fashion Fashion-MNIST
CPU 43.43 44.45 46.45
RTX 4060 (8G) 14.72 13.21 12.43
RTX 3080 (10G) 7.88 7.16 7.17
RTX 4090 (24G) 3.27 3.27 3.27
@@ Conclusion.
We introduced the first gradient-based multiobjective optimization framework called LibMOON, which is implemented based on the PyTorch environment for the research community’s convenience. LibMOON supports mainstream gradient-based MOO methods; the modular design of LibMOON further allows the library to address various MOPs, including discrete Pareto solution optimization, Pareto set learning, multiobjective Bayesian optimization, etc., in a plug-and-play manner. LibMOON can thus be leveraged to quickly yet robustly test new MOO ideas.
Limitations.
The limitations of LibMOON include: (1) the rapid development of gradient-based MOO methods makes it impossible to incorporate all relevant techniques, so some effective methods may be missing; (2) the codebase, primarily based on works from Prof. Qingfu Zhang’s group, may not easily accommodate new methods.
Future Work.
Future work includes: (1) maintaining a user and development community to address issues promptly, and (2) adding newly published or well-known methods to the library as quickly as possible.
@@
LibMOON was made possible with the valuable feedback and comments from a lot of researchers. We appreciate the help from Hongzong Li, Xuehai Pan, Zhe Zhao, Meitong Liu, Weiduo Liao, Weiyu Chen, Baijiong Lin, Dake Bu, Yilu Liu, Song Lai, Cheng Gong, Prof. Jingda Deng, Prof. Ke Shang, Prof. Genghui Li, Prof. Zhichao Lu, and Prof. Tao Qin. We sincerely thank them for their early feedback and contributions.
@@ A.2 Aggregation functions
    $g^{\mathrm{pnorm}}_{\bm{\lambda}}({\bm{\theta}})=\sqrt[\frac{1}{p}]{\left\lVert{\bm{\lambda}}\odot{\bm{L}}({\bm{\theta}})-{\bm{z}}\right\rVert_{p}}.$ (14)
@@
    where $\rho$ is small positive coefficient, usually set as 0.1.
    The contour curve for this functions for a bi-objective case can be found in https://libmoondocs.readthedocs.io/en/latest/gallery/aggfuns.html.
@@ A.3 Metrics
Metrics can be categorized into two groups. The first group evaluates the quality of a set of solutions ${\mathbb{Y}}=\{{\bm{y}}^{(1)},\ldots,{\bm{y}}^{(N)}\}$, with specific metrics such as IGD and FD relying on the known Pareto front for accuracy.
The second group of metrics assesses the quality of individual solutions ${\bm{y}}$ when a preference vector ${\bm{\lambda}}$ is provided.
Group 1: Metrics for a set of solutions
@@
Group 2: Metrics for individual solutions
@@
Those metrics are summarized in Table 13 and also can be found in https://libmoondocs.readthedocs.io/en/latest/apis/Libmoon_metric.html.
@@
PBI Penalty-based intersection The weighted distance between inner product and distance to a preference vector.
@@
A.4 License and usage of LibMOON

### v3

@@ 4 Empirical studies
In this section, we present the empirical results of LibMOON. The experiments were conducted on a personal computer with an i7-10700 CPU and a 10GB RTX3080 GPU. GPU acceleration analysis was performed using RTX3080, 4060, and 4090. The results cover six areas: MOO solvers for synthetic problems, Pareto set learning solvers for synthetic problems, MOO solvers for MTL problems, Pareto set learning solvers for MTL problems, MOBO for synthetic problems, and GPU acceleration (details in Section A.4).
@@ 4.1 MOO solvers for synthetic problems
We report the performance of various finite Pareto solution solvers using the VLMOP2 problem, where $f_{1}({\bm{\theta}})=1-\exp(-\sum_{i=1}^{n}(\theta_{i}-\frac{1}{\sqrt{n}})^{2})$, $f_{2}({\bm{\theta}})=1-\exp(-\sum_{i=1}^{n}(\theta_{i}+\frac{1}{\sqrt{n}})^{2})$. VLMOP2 has been widely studied in the literature [16, 13] since its PF is non-convex  [^4] In MOO, a PF is convex or non-convex based on whether the objective space is convex or non-convex. The PF represents the non-dominated boundary of the objective space.. Some methods (e.g., Agg-LS) fail immediately on this problem since its PF is non-convex. Visualization results are shown by Figure 3 and the numerical results are shown by Table 6. We present the following key findings:
@@
Figure 3: Finite Pareto solutions by ten MOO solvers on VLMOP2 problem.
@@
Method $l_{\mathrm{min}}$ $sl_{\mathrm{min}}$ Spacing Sparsity HV IP Cross Angle PBI
@@
  • All methods were run five times with random seeds; results are presented in the format of (mean)(std).
①. Agg-COSMOS produces rough ‘exact’ Pareto solutions due to a cosine similarity term encouraging Pareto objectives to be close to preference vectors. However, the position of Pareto objectives can not be determined. Agg-LS can only find two endpoints of the PF. ②. Agg-PBI generates “exact” Pareto solutions when the coefficient of $d_{2}$ exceeds a specific value [52]. However, this parameter is challenging to tune, which is influenced by the curvature of a PF. Additionally, PBI may transform a convex multi-objective optimization problem into a non-convex one, making Agg-PBI less recommended. ③. Agg-Tche generates diverse solutions and produces exact Pareto solutions corresponding to the element-wise inverse of the preference vector. Both Agg-Tche and Agg-SmoothTche retain convexity - their aggregation functions keep convex when all objectives are convex. ④. HVGrad updates the decision variable using the hypervolume gradient, resulting in the largest hypervolume. PMTL is a two-stage method. In the first stage, solutions are updated to specific regions and in the second stage, solutions are updated to Pareto solutions constrained in these specific regions. MOO-SVGD’s update direction has two conflicting goals: promoting diversity and ensuring convergence. This conflict makes MOO-SVGD less stable and take more iterations to converge. ⑤. Among these methods, MGDA-UB, Random, Agg-PBI, and MOO-SVGD exhibit relatively large deviations. MGDA-UB and Random generate arbitrary Pareto solutions due to their computation nature. Agg-PBI results in non-convex aggregation functions, leading to variable solutions based on different initializations.
Conclusion. For synthetic problems, the most recommended method is Agg-Tche since (1) it keeps convexity, (2) it finds ‘exact’ Pareto solutions under quite mild conditions (3) it does not need to calculate the Jacobian matrix for each iteration.
@@ 4.2 Pareto set learning on synthetic problems
In this section, we present PSL results (Figures 4 and 7) also using VLMOP2. We have:
@@
Method $l_{\mathrm{min}}$ $sl_{\mathrm{min}}$ Spacing Sparsity HV IP Cross Angle PBI Span
COSMOS 0.045 (0.000) -0.127 (0.000) 1.560 (0.004) 0.525 (0.000) 0.318 (0.000) 0.752 (0.000) 0.950 (0.001) 0.995 (0.000) 0.907 (0.000)
@@
①. PSL with the COSMOS aggregation function fails to find all marginal Pareto solutions because COSMOS does not guarantee the discovery of the entire PS/PF. PSL with linear scalarization function could not fit the two endpoints of the PF. Those PSL results inherit from their base MOO solvers. ②. PSL with the smooth Tchebycheff function finds diverse but non-exact Pareto solutions. In contrast, PSL with Agg-Tche, EPO, and PMGDA as base solvers discovers the entire PS/PF, as all three methods find exact Pareto solutions. By traversing the preference simplex, the model accurately fits the entire PS.
Conclusion. The most recommended method is still Agg-Tche-based PSL since its basic MOO solver Agg-Tche has attractive properties as mentioned in the previous section.
(a) COSMOS
@@ 4.3 MOO solvers for MTL problems
We evaluate the performance of finite Pareto solvers on the Adult dataset, a multitask fairness classification problem. The decision variable ${\bm{\theta}}$ represents the parameters of a fully-connected neural network with $|{\bm{\theta}}|=28033$. The first objective is cross-entropy loss, and the second is the DEO loss [33][Eq. 6]. Key findings from the Figures 5 and 8 are as follows:
①. Agg-LS has two drawbacks: (1) it cannot identify the non-convex part of a PF (as previous section mentioned), and (2) the relationship between preference vectors and Pareto objectives is unknown; different preference vectors may yield duplicate solutions. Agg-PBI and Agg-COSMOS only find a small portion of the PF. ②. Agg-Smooth mTche and Agg-mTche perform well on this task, as they can find (approximate) ‘exact’ Pareto solutions. Once the Pareto front range is known, diverse solutions can be easily found using uniform preference vectors. The Random and MGDA-UB methods only find a single Pareto solution, since the position of this solution cannot be controlled by these methods. ③. Among the three methods that directly find a set of Pareto solutions (MOO-SVGD, PMTL, and HV-Grad), HV-Grad produces the most diverse solutions with the largest hypervolume. PMTL, being a two-stage method, may fail when solutions fall outside the sector due to stochastic factors. MOO-SVGD optimizes both convergence and diversity but is generally unstable based on our tests.
Conclusion. For convex Pareto fronts in MTL problems, Agg-LS is recommended for computational efficiency. However, PMGDA or EPO may offer better convergence and preference-solution correspondence.
@@
Method $l_{\mathrm{min}}$ $sl_{\mathrm{min}}$ Spacing Sparsity HV IP Cross Angle PBI Span
@@
  • All methods were run five times with random seeds; results are presented in the form of (mean)(std).
@@ 4.4 Pareto set learning on MTL
This section presents the Pareto set learning results on the MO-MNIST problem using a hypernet architecture. The Pareto set model was trained for 20 epochs, optimizing approximately 3.24M parameters, with the first and second objectives being the cross-entropy losses for the top-right and bottom-left images. EPO-based or PMGDA-based PSL is not very suitable for this task since manipulating gradient of 3.2M parameters is not efficient. Empirically, for regression tasks, the PF shape is nearly convex. Therefore, Agg-LS is adequate to recover the whole PF. From Figures 6 and 9, we have that Agg-LS significantly outperforms other methods, evidenced by the HV of Agg-LS is much larger than other methods. Furthermore, the training losses across all methods are almost stable after 40 epochs. From the figure, to further reduce the training loss, it needs much more computational resources.
Conclusion. For MTL problems with a convex PF, it is highly recommended to use Agg-LS-based PSL for the sake of computational efficiency.
@@
Method $l_{\mathrm{min}}$ $sl_{\mathrm{min}}$ Spacing Sparsity HV IP Cross Angle PBI Span
COSMOS 0.033 (0.002) -0.152 (0.007) 0.967 (0.202) 0.278 (0.038) 0.512 (0.015) 0.535 (0.028) 9.208 (0.541) 1.221 (0.050) 0.497 (0.043)
@@ Conclusion.
We introduced the first modern gradient-based MOO framework called LibMOON in PyTorch for the research community’s convenience. LibMOON supports more than 20 mainstream gradient-based MOO methods; the modular design of LibMOON further allows the library to address various MOPs via various methods in a plug-and-play manner. LibMOON can thus be leveraged to quickly yet robustly test new MOO ideas.
Limitations
include: (1) rapid developments of gradient-based MOO methods makes it hard to incorporate all methods, so some effective methods may be missing; (2) gradient-based solvers may fail for problems with a number of local optimas.
Future Work
includes (1) maintaining a user and development community to address issues and (2) adding newly published methods as quickly as possible.
@@
LibMOON was made possible with the valuable feedback and comments from a lot of researchers. We appreciate the help from Hongzong Li (for some advice on skipping local optima), Xuehai Pan (for some software design advice), Weiyu Chen (for his code on LoRA PSL), Baijiong Lin (for his advice on automatic code testing), Prof. Hao Wang (for early email communication of HVGrad), Prof. Jingda Deng (for his advice on high-D HV calculation), Prof. Ke Shang (HV computation and RE test suit), Prof. Genghui Li, Prof. Zhichao Lu and Prof. Zhenkun Wang.
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China [GRF Project No. CityU 11215622].
@@ A.2 Aggregation functions
    $g^{\mathrm{pnorm}}_{\bm{\lambda}}({\bm{\theta}})={\left\lVert{\bm{\lambda}}\odot{\bm{L}}({\bm{\theta}})-{\bm{z}}\right\rVert_{p}}.$ (9)
@@
    where $\rho$ is small positive coefficient, which is usually set as 0.1. Contour curves for this functions for a bi-objective case can be found in the LibMOON Doc  [^5] https://libmoondocs.readthedocs.io/en/latest/gallery/aggfuns.html.
@@ A.3 Metrics
Metrics used in LibMOON can be categorized into two groups. The first group evaluates the quality of a set of solutions ${\mathbb{Y}}=\{{\bm{y}}^{(1)},\ldots,{\bm{y}}^{(N)}\}$, with specific metrics such as IGD and FD relying on the known Pareto front for accuracy. The second group of metrics assesses the quality of individual solutions ${\bm{y}}$ when a preference vector ${\bm{\lambda}}$ is provided.
Group 1: Metrics for a set of solutions.
@@
Group 2: Metrics for individual solutions.
@@
Those metrics are summarized in Table 12 and also can be found in the LibMOON document.
@@
PBI Penalty-based intersection The weighted sum distance between inner product and distance to a preference vector.
@@
A.4 GPU acceleration
Table 13: Running time (minutes) for different platforms on MO-MNIST problem.
Platform MO-MNIST MO-Fashion Fashion-MNIST
CPU 43.43 44.45 46.45
RTX 4060 (8G) 14.72 13.21 12.43
RTX 3080 (10G) 7.88 7.16 7.17
RTX 4090 (24G) 3.27 3.27 3.27
A.5 License, usage, and code dependence
@@
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] .
    2. (b)
      Did you describe the limitations of your work? [Yes] , see Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] . LibMOON is a basic optimization library and we do not see direct societal impacts.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [N/A] .
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [Yes] . The assumption is in the theorem itself, e.g., in Theorem 1.
    2. (b)
      Did you include complete proofs of all theoretical results? [No] . This theorem is a restatement of previous results. We have give both the thereom and its proof proper citations.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] . These instruments are provided in the LibMOON Github page: https://github.com/xzhang2523/libmoon.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] . See source code of LibMOON.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] . We have reported standard derivation results.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] . See first paragraph of Section 4.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] . See Section A.5.
    2. (b)
      Did you mention the license of the assets? [Yes] . See Section A.5.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] . Code are provided in https://github.com/xzhang2523/libmoon.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] . LibMOON uses public data such as multiobjective classification and fairness classification.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] .
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] .
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]